OpenStudy (anonymous):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F. F = 9 over 5C + 32 Which of the following formulas correctly solves for C? C = 5F - 32 over 9 C = 5 F over 9 - 32 C = 9 over 5(F – 32) C = 5 over 9(F – 32)

OpenStudy (anonymous):

F = (9/5)C + 32 subtract 32 from both sides F -32 = (9/5)C + 32 -32 simplify F -32 = (9/5)C times both sides by the reciprical of (9/5) (5/9)*(F-32)=(9/5)C*(5/9) simpify (5/9)*(F-32)=C

OpenStudy (anonymous):

So D